# 11footnotetext: co-authors   co-senior authors

### v2

@@ 1 Introduction
Scientific progress is cumulative and open-ended, with each breakthrough standing on the shoulders of countless prior insights. In the same way, our most advanced AI systems are built upon a long lineage of innovations. For instance, transformers (Vaswani et al., 2017), the backbone of current large language models (LLMs) (Brown et al., 2020), did not emerge in isolation but were built upon years of past innovations, such as recurrent neural networks (Linnainmaa, 1970; Amari, 1972; Hopfield, 1982; Rumelhart et al., 1985) and attention mechanisms (Schmidhuber & Huber, 1990; Bahdanau et al., 2015; Kim et al., 2017; Parikh et al., 2016). However, most of today’s AI systems remain bound by fixed, human-designed architectures that learn within predefined boundaries, without the capacity to autonomously rewrite their own source code to self-improve. As a result, each advancement in AI development still leans heavily on human interventions, tethering the pace of progress. This paper investigates the intriguing possibility of safely automating the search for ever-better AI. One can imagine an AI system that, like scientific discovery itself, becomes an engine of its own advancement: building upon its past, recursively improving, and propelling itself toward more advanced capabilities.
@@ 2 Related Work
Open-Endedness. A grand challenge for driving unbounded innovation is designing open-ended AI systems that continuously generate novel and learnable artifacts (Stanley et al., 2017). Hughes et al. (2024) characterized open-endedness as a system’s capacity to generate sequences of artifacts that are both novel and learnable from an observer’s perspective. A central difficulty lies in structuring and exploring vast search spaces to consistently produce artifacts that are interesting to humans (Clune, 2019; Jiang et al., 2023). Early progress drew on quality-diversity algorithms, goal-directed exploration, intrinsic motivation, and learning-progress frameworks (Pugh et al., 2016; Ecoffet et al., 2019; Lehman & Stanley, 2011; Oudeyer et al., 2007), while recent advances leverage large-scale foundation models (FMs) as proxies for human interestingness and versatile engines for generating and evaluating novel behaviors across diverse domains (Brown et al., 2020; Hu et al., 2025; Zhang et al., 2024b). However, these approaches have yet to close the self-referential self-improvement loop, meaning improvements on downstream tasks do not translate into enhanced capabilities for self-modification or the acceleration of further innovations. We aim to mimic the acceleration of science and technology, where new tools and discoveries catalyze the creation of even more discoveries. How can we emulate nature’s arc of evolution, which bends not only toward complexity but also an ever greater capacity to evolve (Dawkins, 2019; Gerhart & Kirschner, 2007; Hendrikse et al., 2007)?
@@
Meta-Learning FM Agents. Many FM-based agents are handcrafted. Some building blocks include prompt engineering (Chen et al., 2023; Schulhoff et al., 2024), chain-of-thought (Wei et al., 2022; Yao et al., 2023; Hu & Clune, 2024; Guo et al., 2025; Lightman et al., 2023; Muennighoff et al., 2025; Zelikman et al., 2024a), self-reflection (Shinn et al., 2023; Yao et al., 2023; Madaan et al., 2023), multi-agent debate (Zhuge et al., 2023; Liang et al., 2023; Khan et al., 2024), memory (Liu et al., 2023; Zhong et al., 2024; Modarressi et al., 2023), temperature sampling (Zhu et al., 2024), and retrieval augmented generation (Lewis et al., 2020). The manual composition of these components limits the system’s abilities to the ingenuity of its human designer. More recently, several meta-learning approaches have emerged that leverage FM to automatically optimize prompts (Fernando et al., 2024; † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR; Khattab et al., 2023; Cheng et al., 2024; Yuksekgonul et al., 2024; Yuan et al., 2024) and design agentic modules (Zhang et al., 2024c; Zhou et al., 2024; Yin et al., 2024; Zhuge et al., 2024; Rosser & Foerster, 2025; Zhang et al., 2025a; Ye et al., 2025; Gao et al., 2025; Nie et al., 2025; Su et al., 2025; Zhang et al., 2025b; Niu et al., 2025). The Automated Design of Agentic Systems (ADAS, Hu et al., 2025) iteratively generates downstream agents with a fixed meta-agent, evaluates them against a target benchmark, and incorporates feedback to refine subsequent generations. In contrast, the DGM is a single system that both solves downstream tasks (i.e., coding problems) and refines its own implementation (i.e., its codebase), removing the need for a fixed, handcrafted meta-agent and enabling self-referential improvements.
@@
Self-Improving AI. Early on, various researchers outlined theoretical and conceptual approaches to self-improvement (Good, 1966; Schmidhuber, 1987; 2007). Some practical approaches to automated self-improvement include systems defined by neural network weight parameterizations (Schmidhuber, 1993; Hall, 2007; Hobbhahn, 2025; Kirsch & Schmidhuber, 2022; Irie et al., 2022; 2025; Lu et al., 2023; Havrilla et al., 2024b). Metz et al. (2021) developed a gradient-based optimizer that is self-referentially meta-trained using a variant of population-based training (Jaderberg et al., 2017). Lange et al. (2023) extended this approach to gradient-free learning. Silver et al. (2017) used self-play to continuously evolve agents, achieving superhuman performance in challenging domains such as chess and Go. More closely related to the DGM are recent approaches that leverage FM-based agents for self-improvement (Yin et al., 2024; Robeyns et al., 2025; Hu et al., 2024; Zelikman et al., 2024b; Huang et al., 2022; Singh et al., 2023). Zelikman et al. (2024b) use a meta-agent to generate downstream agents, updating the meta-agent based on the meta-utility derived from the generated solutions. Yin et al. (2024) use a single system to both solve downstream tasks and recursively modify itself. However, the downstream tasks or the meta-utility do not always align with the capabilities required for self-improvement. In the DGM, improvement in downstream tasks directly reflects an increase in self-improvement ability, enabling the potential for self-accelerating progress. Most similar is concurrent work by Robeyns et al. (2025), which also has a single agent recursively solving coding problems and modifying its own codebase. The main difference from Robeyns et al. (2025) is that the DGM has an open-ended exploration loop, encouraging self-modifications beyond immediate performance gains and thus avoiding stagnation in suboptimal states. Appendix˜B also discusses additional related work on program synthesis and Darwinian evolution.
@@ 3 Darwin Gödel Machine
A Gödel Machine is a theoretical idea of an AI that searches for ways that provably improve itself (Schmidhuber, 2007). In this paper, we propose Darwin Gödel Machine (DGM), an attempt to realize the long-held dream of creating a Gödel Machine. The DGM relaxes the Gödel Machine’s impractical requirement of theoretically proving that a change will improve the system, instead requiring empirical evidence from experiments to demonstrate that a proposed new version enhances performance. Additionally, since the DGM relies on empirical evidence of improvement, it may get stuck in a local optimum within the vast search space of possible systems (i.e., all computable algorithms). To address this, the DGM maintains an archive of discovered solutions during the search, facilitating open-ended exploration rather than relying on evolving a single solution. Since the principles echo Darwinian evolution (Darwin, 2023) (Appendix˜B), where new innovations emerge by selecting an entity from an archive of previously discovered solutions, modifying it, and keeping it if it is interestingly new (Zhang et al., 2024b; Faldor et al., 2025; Stanley & Lehman, 2015), we call our algorithm a Darwin Gödel Machine (Figure˜1).
@@
Population-based Open-ended Exploration. Starting from a single initial coding agent, the DGM builds an archive of all discovered agents. In each iteration, the DGM selects parent agents to self-modify and branch off to produce new agents. Parent selection is roughly proportional to each agent’s performance score and the number of its children with codebase-editing functionality (Section˜C.2). Each selected parent analyzes its own benchmark evaluation logs, proposes the next feature to implement, and receives this proposal as a problem statement to execute (Section˜C.3). The parent then implements the suggested feature into its own codebase, generating a new coding agent. Each newly generated agent is quantitatively evaluated on a chosen coding benchmark to estimate its coding abilities. Only agents that compile successfully and retain the ability to edit a given codebase are added to the DGM archive, as only they can continue self-modification. All others are discarded. The cycle of parent selection, self-modification, and evaluation continues, progressively growing the archive of solutions. Importantly, we note that archived solutions can serve as stepping stones that result in improvements much later than their original discovery, making our approach substantially different from hill-climbing agentic design approaches (Robeyns et al., 2025). Currently, the open-ended exploration process (i.e., archive maintenance, parent selection) is fixed and not modifiable by the DGM, which we leave as an avenue for future work. Section˜C.4 shows the pseudocode for the DGM algorithm.
@@ 4.1 Experiment Setup
The DGM is initialized with a single coding agent. This base agent is built around a frozen FM and augmented with tool use capabilities (Schick et al., 2023; Anthropic, 2024b). The FM autonomously determines its action at each step, using whichever tool it deems appropriate. The base agent has access to two tools: a Bash tool for executing bash commands, and an edit tool for viewing and editing directories and entire files (Section˜C.1). The agent receives a single input prompt that specifies the location of the target repository, the problem statement, and the repository’s testing framework (Section˜C.1). We run the DGM for 80 iterations (generating one new agent per iteration), with two iterations running in parallel for SWE-bench and four for Polyglot (Section˜C.2). During self-modifications, coding agents are powered by Claude 3.5 Sonnet (New) (Anthropic, 2024a) in both SWE-bench and Polyglot experiments. During benchmark evaluation, coding agents are powered by Claude 3.5 Sonnet (New) for SWE-bench and o3-mini (OpenAI, 2025) for Polyglot (Section˜D.1).
@@ 4.2 Benchmarks
Both SWE-bench and Polyglot are benchmarks that evaluate how well AI agents solve coding tasks automatically. Given a code repository and a task instruction, the agent is expected to make changes to the repository in order to fulfill the task. Both SWE-bench and Polyglot are widely used benchmarks (Zhang et al., 2024d; a; Xia et al., 2024; Cao et al., 2024; Google DeepMind, 2025; Gauthier, 2024) that require the AI agent to navigate a code repository, understand the interplay between functions in different files, and spot small errors in convoluted code. SWE-bench only has Python tasks, while Polyglot has tasks in multiple programming languages. Another difference is that each SWE-bench task may require edits to multiple files, whereas each Polyglot task primarily involves implementing a solution from scratch in a single file (although the agent still needs to examine other files to understand what changes are necessary), resulting in fewer file edits overall. We evaluate the DGM on two popular benchmarks that assess different aspects of coding tasks to validate the algorithm’s effectiveness across various use cases.
@@ 4.3 Baselines
To isolate and evaluate the contribution of the DGM’s core components, we compare the DGM against two baselines: DGM without self-improving agents (DGM w/o self-improve) and DGM without open-ended exploration (DGM w/o open-ended exploration). DGM w/o self-improve replicates the approach of ADAS (Hu et al., 2025) in this setting, wherein the meta agent responsible for modifying the coding agents remains fixed as the base agent throughout the experiment (Algorithm˜2). This baseline allows us to assess the role of iterative self-improvement in accelerating the evolution of better coding agents. DGM w/o open-ended exploration eliminates the use of an archive and always self-modifies the latest stored version of itself (Algorithm˜3). If a coding agent self-modifies to the point where it loses the basic functionality required to edit a codebase, it can no longer modify itself or solve any coding task. Therefore, DGM w/o open-ended exploration retains the latest version of itself that still maintains the basic functionality for codebase editing. This baseline allows us to evaluate the impact of having an archive and the well-documented beneficial principles of open-ended exploration (Clune, 2019; Stanley & Lehman, 2015; Zhang et al., 2024b; Fernando et al., 2024; Lee et al., 2020; Samvelyan et al., 2024; Colas et al., 2022b) in guiding the agent’s evolution.
@@ 5 Safety Discussion
Conversely, a significant potential benefit of the self-improvement paradigm is that it could, in principle, be directed toward enhancing safety and interpretability themselves. We conduct a preliminary investigation into how the DGM can be deployed in AI safety settings to develop countermeasures for FM hallucination (Appendix˜H). Just as the DGM learns to improve its coding capabilities, it could potentially discover and integrate better internal safeguards or modify itself for greater transparency (e.g., incorporating principles akin to Constitutional AI (Bai et al., 2022)), if such properties were included in its evaluation criteria (Rosser & Foerster, 2025). This suggests a promising, albeit challenging, pathway in which self-improvement becomes a tool for building more trustworthy AI systems. Additional research could also explore weaving Constitutional AI in from the start, though the challenge would be incentivizing the system to retain these directives (an option worth exploring is to create an unmodifiable part of the system to be able to evaluate at halt the rest).
The DGM demonstrates the potential of self-improving AI while still operating within safe research boundaries due to the current limitations of frontier FMs and effective mitigations like sandboxing. We include this safety discussion proactively to raise awareness about the emerging prospect of self-improving AI systems and their associated safety implications, particularly as these systems inevitably become more capable (Yudkowsky et al., 2008; Bostrom, 2002; Ecoffet et al., 2020; Bengio et al., 2024; Clune, 2019). Accordingly, we advocate for continued investigation into the safe and beneficial evolution of AI-Generating Algorithms (Clune, 2019) and self-improving systems.

### v3

# 11footnotetext: co-first authors   co-senior authors

@@ 1 Introduction
Scientific progress is cumulative and open-ended, with each breakthrough standing on the shoulders of countless prior insights. In the same way, our most advanced AI systems are built upon a long lineage of innovations. For instance, transformers (Vaswani et al., 2017), the backbone of current large language models (LLMs) (Brown et al., 2020), did not emerge in isolation but were built upon years of past innovations, such as recurrent neural networks (Linnainmaa, 1970; Amari, 1972; Hopfield, 1982; Rumelhart et al., 1985) and attention mechanisms (Schmidhuber and Huber, 1990; Bahdanau et al., 2015; Kim et al., 2017; Parikh et al., 2016). However, most of today’s AI systems remain bound by fixed, human-designed architectures that learn within predefined boundaries, without the capacity to autonomously rewrite their own source code to self-improve. As a result, each advancement in AI development still leans heavily on human interventions, tethering the pace of progress. This paper investigates the intriguing possibility of safely automating the search for ever-better AI. One can imagine an AI system that, like scientific discovery itself, becomes an engine of its own advancement: building upon its past, recursively improving, and propelling itself toward more advanced capabilities.
@@ 2 Related Work
Open-Endedness. A grand challenge for driving unbounded innovation is designing open-ended AI systems that continuously generate novel and learnable artifacts (Stanley et al., 2017). Hughes et al. (2024) characterized open-endedness as a system’s capacity to generate sequences of artifacts that are both novel and learnable from an observer’s perspective. A central difficulty lies in structuring and exploring vast search spaces to consistently produce artifacts that are interesting to humans (Clune, 2019; Jiang et al., 2023). Early progress drew on quality-diversity algorithms, goal-directed exploration, intrinsic motivation, and learning-progress frameworks (Pugh et al., 2016; Ecoffet et al., 2019; Lehman and Stanley, 2011; Oudeyer et al., 2007), while recent advances leverage large-scale foundation models (FMs) as proxies for human interestingness and versatile engines for generating and evaluating novel behaviors across diverse domains (Brown et al., 2020; Hu et al., 2025; Zhang et al., 2024b). However, these approaches have yet to close the self-referential self-improvement loop, meaning improvements on downstream tasks do not translate into enhanced capabilities for self-modification or the acceleration of further innovations. We aim to mimic the acceleration of science and technology, where new tools and discoveries catalyze the creation of even more discoveries. How can we emulate nature’s arc of evolution, which bends not only toward complexity but also an ever greater capacity to evolve (Dawkins, 2019; Gerhart and Kirschner, 2007; Hendrikse et al., 2007)?
@@
Meta-Learning FM Agents. Many FM-based agents are handcrafted. Some building blocks include prompt engineering (Chen et al., 2023; Schulhoff et al., 2024), chain-of-thought (Wei et al., 2022; Yao et al., 2023; Hu and Clune, 2024; Guo et al., 2025; Lightman et al., 2023; Muennighoff et al., 2025; Zelikman et al., 2024a), self-reflection (Shinn et al., 2023; Yao et al., 2023; Madaan et al., 2023), multi-agent debate (Zhuge et al., 2023; Liang et al., 2023; Khan et al., 2024), memory (Liu et al., 2023; Zhong et al., 2024; Modarressi et al., 2023), temperature sampling (Zhu et al., 2024), and retrieval augmented generation (Lewis et al., 2020). The manual composition of these components limits the system’s abilities to the ingenuity of its human designer. More recently, several meta-learning approaches have emerged that leverage FM to automatically optimize prompts (Fernando et al., 2024; (FAIR)† et al., 2022; Khattab et al., 2023; Cheng et al., 2024; Yuksekgonul et al., 2024; Yuan et al., 2024) and design agentic modules (Zhang et al., 2024c; Zhou et al., 2024; Yin et al., 2024; Zhuge et al., 2024; Rosser and Foerster, 2025; Zhang et al., 2025a; Ye et al., 2025; Gao et al., 2025; Nie et al., 2025; Su et al., 2025; Zhang et al., 2025b; Niu et al., 2025). The Automated Design of Agentic Systems (ADAS, Hu et al., 2025) iteratively generates downstream agents with a fixed meta-agent, evaluates them against a target benchmark, and incorporates feedback to refine subsequent generations. In contrast, the DGM is a single system that both solves downstream tasks (i.e., coding problems) and refines its own implementation (i.e., its codebase), removing the need for a fixed, handcrafted meta-agent and enabling self-referential improvements.
@@
Self-Improving AI. Early on, various researchers outlined theoretical and conceptual approaches to self-improvement (Good, 1966; Schmidhuber, 1987; 2007). Some practical approaches to automated self-improvement include systems defined by neural network weight parameterizations (Schmidhuber, 1993; Hall, 2007; Hobbhahn, 2025; Kirsch and Schmidhuber, 2022; Irie et al., 2022; 2025; Lu et al., 2023; Havrilla et al., 2024b). Metz et al. (2021) developed a gradient-based optimizer that is self-referentially meta-trained using a variant of population-based training (Jaderberg et al., 2017). Lange et al. (2023) extended this approach to gradient-free learning. Silver et al. (2017) used self-play to continuously evolve agents, achieving superhuman performance in challenging domains such as chess and Go. More closely related to the DGM are recent approaches that leverage FM-based agents for self-improvement (Yin et al., 2024; Robeyns et al., 2025; Hu et al., 2024; Zelikman et al., 2024b; Huang et al., 2022; Singh et al., 2023). Zelikman et al. (2024b) use a meta-agent to generate downstream agents, updating the meta-agent based on the meta-utility derived from the generated solutions. Yin et al. (2024) use a single system to both solve downstream tasks and recursively modify itself. However, the downstream tasks or the meta-utility do not always align with the capabilities required for self-improvement. In the DGM, improvement in downstream tasks directly reflects an increase in self-improvement ability, enabling the potential for self-accelerating progress. Most similar is concurrent work by Robeyns et al. (2025), which also has a single agent recursively solving coding problems and modifying its own codebase. The main difference from Robeyns et al. (2025) (and also Zelikman et al. (2024b); Yin et al. (2024)) is that the DGM has an open-ended exploration loop, encouraging self-modifications beyond immediate performance gains and thus avoiding stagnation in suboptimal states. Appendix˜B also discusses additional related work on program synthesis and Darwinian evolution.
@@ 3 Darwin Gödel Machine
A Gödel Machine is a theoretical idea of an AI that searches for ways that provably improve itself (Schmidhuber, 2007). In this paper, we propose Darwin Gödel Machine (DGM), an attempt to realize the long-held dream of creating a Gödel Machine. The DGM relaxes the Gödel Machine’s impractical requirement of theoretically proving that a change will improve the system, instead requiring empirical evidence from experiments to demonstrate that a proposed new version enhances performance. Additionally, since the DGM relies on empirical evidence of improvement, it may get stuck in a local optimum within the vast search space of possible systems (i.e., all computable algorithms). To address this, the DGM maintains an archive of discovered solutions during the search, facilitating open-ended exploration rather than relying on evolving a single solution. Since the principles echo Darwinian evolution (Darwin, 2023) (Appendix˜B), where new innovations emerge by selecting an entity from an archive of previously discovered solutions, modifying it, and keeping it if it is interestingly new (Zhang et al., 2024b; Faldor et al., 2025; Stanley and Lehman, 2015), we call our algorithm a Darwin Gödel Machine (Figure˜1).
@@
Population-based Open-ended Exploration. Starting from a single initial coding agent, the DGM builds an archive of all discovered agents. In each iteration, the DGM selects parent agents to self-modify and branch off to produce new agents. Parent selection is roughly proportional to each agent’s performance score and inversely proportional to the number of its children with codebase-editing functionality (Section˜C.2). This favors high-performing agents that have been underexplored (i.e., have fewer existing children), thereby promoting both exploitation of strong performers and exploration of promising but less-sampled lineages. All agents retain a non-zero selection probability, ensuring that any path to improvement remains feasible given sufficient compute. Each selected parent analyzes its own benchmark evaluation logs, proposes the next feature to implement, and receives this proposal as a problem statement to execute (Section˜C.3). The parent then implements the suggested feature into its own codebase, generating a new coding agent. Each newly generated agent is quantitatively evaluated on a chosen coding benchmark to estimate its coding abilities. Only agents that compile successfully and retain the ability to edit a given codebase are added to the DGM archive, as only they can continue self-modification. All others are discarded. The cycle of parent selection, self-modification, and evaluation continues, progressively growing the archive of solutions. Importantly, we note that archived solutions can serve as stepping stones that result in improvements much later than their original discovery, making our approach substantially different from hill-climbing agentic design approaches (Robeyns et al., 2025). Currently, the open-ended exploration process (i.e., archive maintenance, parent selection) is fixed and not modifiable by the DGM, which we leave as an avenue for future work. Section˜C.4 shows the pseudocode for the DGM algorithm.
@@ 4.1 Experiment Setup
The DGM is initialized with a single coding agent. This base agent is built around a frozen FM and augmented with tool use capabilities (Schick et al., 2023; Anthropic, 2024b). The FM autonomously determines its action at each step, using whichever tool it deems appropriate. The base agent has access to two tools: a Bash tool for executing bash commands, and an edit tool for viewing and editing directories and entire files (Section˜C.1). The agent receives a single input prompt that specifies the location of the target repository, the problem statement, and the repository’s testing framework (Section˜C.1). The base agent is intentionally lightweight to provide a simple starting point from which self-improvement can be clearly demonstrated. We run the DGM for 80 iterations (generating one new agent per iteration), with two iterations running in parallel for SWE-bench and four for Polyglot (Section˜C.2). During self-modifications, coding agents are powered by Claude 3.5 Sonnet (New) (Anthropic, 2024a) in both SWE-bench and Polyglot experiments. During benchmark evaluation, coding agents are powered by Claude 3.5 Sonnet (New) for SWE-bench and o3-mini (OpenAI, 2025) for Polyglot (Section˜D.1).
@@ 4.2 Benchmarks
We evaluate the DGM on two popular benchmarks that assess different aspects of coding tasks to validate the algorithm’s effectiveness across various use cases. Both SWE-bench and Polyglot are benchmarks that evaluate how well AI agents solve coding tasks automatically. Given a code repository and a task instruction, the agent is expected to make changes to the repository in order to fulfill the task. Both SWE-bench and Polyglot are widely used benchmarks (Zhang et al., 2024d; a; Xia et al., 2024; Cao et al., 2024; Google DeepMind, 2025; Gauthier, 2024) that require the AI agent to navigate a code repository, understand the interplay between functions in different files, and spot small errors in convoluted code. SWE-bench only has Python tasks, while Polyglot has tasks in multiple programming languages. Another difference is that each SWE-bench task may require edits to multiple files, whereas each Polyglot task primarily involves implementing a solution from scratch in a single file (although the agent still needs to examine other files to understand what changes are necessary), resulting in fewer file edits overall.
@@ 4.3 Baselines
To isolate and evaluate the contribution of the DGM’s core components, we compare the DGM against two baselines: DGM without self-improving agents (DGM w/o self-improve) and DGM without open-ended exploration (DGM w/o open-ended exploration). DGM w/o self-improve replicates the approach of ADAS (Hu et al., 2025) in this setting, wherein the meta agent responsible for modifying the coding agents remains fixed as the base agent throughout the experiment (Algorithm˜2). This baseline allows us to assess the role of iterative self-improvement in accelerating the evolution of better coding agents. DGM w/o open-ended exploration eliminates the use of an archive and always self-modifies the latest stored version of itself (Algorithm˜3). If a coding agent self-modifies to the point where it loses the basic functionality required to edit a codebase, it can no longer modify itself or solve any coding task. Therefore, DGM w/o open-ended exploration retains the latest version of itself that still maintains the basic functionality for codebase editing. This baseline allows us to evaluate the impact of having an archive and the well-documented beneficial principles of open-ended exploration (Clune, 2019; Stanley and Lehman, 2015; Zhang et al., 2024b; Fernando et al., 2024; Lee et al., 2020; Samvelyan et al., 2024; Colas et al., 2022b) in guiding the agent’s evolution.
@@ 5 Safety Discussion
Conversely, a significant potential benefit of the self-improvement paradigm is that it could, in principle, be directed toward enhancing safety and interpretability themselves. We conduct a preliminary investigation into how the DGM can be deployed in AI safety settings to develop countermeasures for FM hallucination (Appendix˜H). Just as the DGM learns to improve its coding capabilities, it could potentially discover and integrate better internal safeguards or modify itself for greater transparency (e.g., incorporating principles akin to Constitutional AI (Bai et al., 2022)), if such properties were included in its evaluation criteria (Rosser and Foerster, 2025). This suggests a promising, albeit challenging, pathway in which self-improvement becomes a tool for building more trustworthy AI systems. Additional research could also explore weaving Constitutional AI in from the start, though the challenge would be incentivizing the system to retain these directives (an option worth exploring is to create an unmodifiable part of the system to be able to evaluate at halt the rest).
The DGM demonstrates the potential of self-improving AI while still operating within safe research boundaries due to the current limitations of frontier FMs and effective mitigations like sandboxing. Appendix˜I presents additional discussion on broader safety uncertainties. We include this safety discussion proactively to raise awareness about the emerging prospect of self-improving AI systems and their associated safety implications, particularly as these systems inevitably become more capable (Yudkowsky and others, 2008; Bostrom, 2002; Ecoffet et al., 2020; Bengio et al., 2024; Clune, 2019). Accordingly, we advocate for continued investigation into the safe and beneficial evolution of AI-Generating Algorithms (Clune, 2019) and self-improving systems.